# Diffusion models for multivariate subsurface generation and efficient probabilistic inversion

### v1

@@ Abstract
Diffusion models offer stable training and state-of-the-art performance for deep generative modeling tasks. Here, we consider their use in the context of multivariate subsurface modeling and probabilistic inversion. We first demonstrate that diffusion models enhance multivariate modeling capabilities compared to variational autoencoders and generative adversarial networks. In diffusion modeling, the generative process involves a comparatively large number of time steps with update rules that can be modified to account for conditioning data. We propose different corrections to the popular Diffusion Posterior Sampling approach by Chung et al. (2023). In particular, we introduce a likelihood approximation accounting for the noise-contamination that is inherent in diffusion modeling. We assess performance in a multivariate geological scenario involving facies and correlated acoustic impedance. Conditional modeling is demonstrated using both local hard data (well logs) and nonlinear geophysics (fullstack seismic data). Our tests show significantly improved statistical robustness, enhanced sampling of the posterior probability density function and reduced computational costs, compared to the original approach. The method can be used with both hard and indirect conditioning data, individually or simultaneously. As the inversion is included within the diffusion process, it is faster than other methods requiring an outer-loop around the generative model, such as Markov chain Monte Carlo.
@@ 1 Introduction
Predicting the spatial distribution of subsurface properties is a key challenge in the Earth and environmental sciences. Probabilistic inverse modeling (Grana et al., 2021; Rubin and Hubbard, 2005) aims to approximate the posterior probability density function (pdf) by integrating prior pdfs with observed data in the form of a likelihood function (Tarantola, 2005). Conditioning data may consist of local measurements (e.g., well logs) or indirect geophysical data describing averaged subsurface properties through linear or nonlinear physical forward operators. Prior pdfs expressed as explicit statistical models (e.g., two-point geostatistical models) are computationally convenient, yet are often gross simplifications of geological heterogeneity, leading to biased predictions and incomplete uncertainty quantification (e.g., Linde et al., 2017; Gómez-Hernández and Wen, 1998; Neuweiler et al., 2011; Journel and Deutsch, 1993; Zhou et al., 2018). Alternatively, prior geological knowledge can be represented using 2D or 3D gridded data, defined as training images (TI), which implicitly describe lower- and higher-order spatial relationships. Multiple-Point Statistics (MPS) simulation methods (Guardiano and Srivastava, 1993; Strebelle, 2002; Mariethoz et al., 2010a; Mariethoz and Caers, 2014) rely on TIs to draw geologically-realistic model realizations and have found wide applicability in both hydro-geology (e.g., Høyer et al., 2017; Le Coz et al., 2017; Straubhaar et al., 2020) and reservoir characterization (e.g., Melnikova et al., 2015; Caers and Zhang, 2004, & references therein). These realizations can easily be made to honor local observations (Straubhaar et al., 2016; Hansen et al., 2018; Straubhaar and Renard, 2021), but struggle with high computational costs and poor posterior exploration when used in the context of nonlinear inversion (e.g., Zahner et al., 2016; Levy et al., 2024; Laloy et al., 2016; Hansen et al., 2012; Mariethoz et al., 2010b).
@@
The substantial progress of deep generative models over the past decade (Goodfellow et al., 2016) have driven significant interest to leverage such techniques in geoscience. Initial work considered convolution-based deep neural networks, such as variational autoencoders (VAE; Kingma and Welling, 2013) and generative adversarial networks (GAN; Goodfellow et al., 2014). These methods can encode complex geological patterns of a TI through a low-dimensional manifold, defining prior pdf models (Laloy et al., 2017, 2018). Therefore, inversion can be performed within the latent space of such pre-trained networks (Laloy et al., 2017, 2018; Oliviero-Durmus et al., 2025) using sampling methods, such as Markov chain Monte Carlo (MCMC; Laloy et al., 2017, 2018; Levy et al., 2021; Mosser et al., 2020; Liu et al., 2022) and Sequential Monte Carlo (SMC; Doucet et al., 2001; Amaya et al., 2021), ensemble-based approaches (Canchumuni et al., 2019; Liu and Grana, 2020; Mo et al., 2020), deterministic optimization (Laloy et al., 2019; Dupont et al., 2018), or variational inference (Chan and Elsheikh, 2019; Levy et al., 2023; Miele et al., 2024). Contrarily to MPS algorithms, conditioning to local data remains cumbersome. One can train the generative models as inference methods on specific conditional data, but this necessitates retraining for each application task (e.g. Mosser et al., 2018; Zhang et al., 2021; Laloy et al., 2020; Feng et al., 2024; Miele and Azevedo, 2024). Moreover, GANs are prone to unstable training and mode collapse (e.g., Lucic et al., 2018), and the non-linearity of the generative process often adversely affects the performance of the inversion (Lopez-Alvis et al., 2021; Laloy et al., 2019). Formulations based on VAEs are more stable, but offer rather lossy representations (e.g., Levy et al., 2023; Miele et al., 2024).
Diffusion models (DM; Ho et al., 2020; Sohl-Dickstein et al., 2015) represent the current state-of-the-art in deep generative modeling, outperforming the accuracy of GANs without being affected by mode collapse or unstable training (e.g., Dhariwal and Nichol, 2021). These methods are based on reversing a data corruption process transforming samples of a target distribution (the TI) into those of a known distribution (typically Gaussian). The generative process involves iteratively denoising samples originating from this distribution, with the denoising trajectory being learned by a deep neural network (e.g., Song et al., 2021a; Ho et al., 2020; Song et al., 2021b; Karras et al., 2022). So far, relatively few works studied DMs in the context of geological subsurface modeling. For example, Mosser (2023) highlighted that DM, in an autoregressive formulation (Hoogeboom et al., 2022), share similarities to sequential geostatistical modeling, and demonstrated promising multi-scale performance in modeling 2D realizations of binary fluvial channel distributions. Xu et al. (2024) demonstrated accuracy of DM in an analogous geological scenario, and proposed a conditioning approach, where the DM is trained with local point observations incorporated as masked sparse data in additional input channels. Diffusion models were also successfully applied by Aouf et al. (2025) for 2D and 3D modeling of complex porous geological media. The authors further performed conditional modeling with respect to total porosity, by training a DM on embedded representation of the conditioning data. Lee et al. (2025) considered a similar conditional approach with latent diffusion models (LDM; Rombach et al., 2022), to generate multi-facies realizations conditioned on well log data. Here, the diffusion and its conditioning are performed within the latent space of a pre-trained encoder-decoder network. Ovanger et al. (2025) extend the work of Lee et al. (2025) by providing a thorough geostatistical assessment of LDM against truncated Gaussian random field modeling (Matheron et al., 1987; Mannseth, 2014). The authors observed good performance for most of the metrics considered, but note significant biases limiting a full prior representation and errors in the conditional realizations. Finally, Di Federico and Durlofsky (2025) train an unconditioned LDM for facies modeling and combine it with an ensemble-based method for data assimilation in the context of history matching. To the best of our knowledge, the latter is the only study to date dealing with subsurface inverse problems using DMs.
@@
Inverse modeling with DMs is of broad interest in computer vision and biomedical imaging; we refer to Daras et al. (2024) and Zhao et al. (2025) for comprehensive overviews. Analogously to GANs and VAEs, direct training on conditioning data is a popular approach (e.g., Ho and Salimans, 2021; Dhariwal and Nichol, 2021; Whang et al., 2022; Liu et al., 2023; Xu et al., 2024; Lee et al., 2025; Aouf et al., 2025), but implies computationally demanding re-training when considering different data types or designs. Alternatively, one can rely on inference frameworks that iterate outside the generative process of pre-trained unconditional DMs. Conventional sampling algorithms (e.g., MCMC) are often inefficient in this setting, due to the DM’s large dimensionality and comparatively long generation times (Zhao et al., 2025); more efficient variational inference algorithms have been proposed for posterior approximation with surrogate modeling, but they often struggle to capture complex or multimodal posterior distributions (e.g., Mardani et al., 2023; Alkan et al., 2023; Feng and Bouman, 2024). Ensemble-based approaches (e.g., Di Federico and Durlofsky, 2025) share similar issues of comparatively high computational costs. Even if LDMs can reduce dimensionality, the associated encoder-decoder networks (e.g., VAE or GAN) introduce additional complications such as reduced modeling ability (e.g., bias or smoothing; Ovanger et al., 2025) and non-linearities (Daras et al., 2024).
@@
Another type of conditioning approach leverages the sequential nature of the denoising process. For example, SMC can be used to progressively sample, at different denoising steps, a posterior distribution that evolves towards the posterior pdf (e.g., Trippe et al., 2022; Cardoso et al., 2023; Wu et al., 2024; Dou and Song, 2024). These asymptotically exact methods work under the assumption of infinite sampling and linear physics. Moreover, the denoising process itself can be adapted to achieve approximate Bayesian posterior sampling at lower computational cost (Song et al., 2021b), provided that the intractable likelihood is approximated at each denoising step. The approximated likelihood score is then combined with the pre-trained denoising score to steer the denoising generative process towards posterior samples. Among numerous such approaches (see e.g., Daras et al., 2024), Diffusion Posterior Sampling (DPS) (Chung et al., 2023) has gained wide popularity as it allows inference for both linear and nonlinear inverse problems. Nonetheless, this class of conditional samplers are affected by their likelihood approximation considered (Daras et al., 2024). For instance, the accuracy of DPS decreases with decreasing conditioning data noise magnitude and increasing nonlinearity of the problem at hand. Therefore, various adjustments to the theoretical framework are used in the implementation of DPS, including ad-hoc weighting of the likelihood score. Several studies have highlighted these limitations and propose partial improvements for linear inverse problems (e.g., Hamidi and Yang, 2025; Boys et al., 2024; Yismaw et al., 2025; Song et al., 2023).
@@
In this work, we consider DMs for multivariate subsurface geological modeling and inversion tasks, proposing corrections to the DPS approach. The improved likelihood approximation of our corrected DPS (CDPS) accounts for the model’s denoising uncertainties; these are propagated into the data domain for likelihood evaluation, accounting for linearized Gaussian approximations when considering nonlinear geophysics conditioning cases, following Friedli and Linde (2024) and Linde et al. (2017). The modified formulation removes the need to arbitrarily weight the likelihood score, and aims at increased robustness. We further highlight inconsistencies in the original implementation of DPS by Chung et al. (2023) implemented with Denoising Diffusion Probabilistic and Implicit Modeling (DDPM and DDIM; Ho et al., 2020; Song et al., 2021a) and suggest corrections. For our study, we use a state-of-the-art DM implementation (EDM; Karras et al., 2022) to model a bivariate prior pdf describing sedimentary channel sequences and correlated acoustic impedances. We first demonstrate substantial increases in prior modeling performance against a GAN and a VAE previously introduced for analogous tasks (Miele et al., 2024). Then, we use CDPS to condition realizations on both local observations (well logs) and nonlinear geophysical data (fullstack seismic data), considering Gaussian data noise of different magnitudes. We demonstrate that CDPS yields significantly improved accuracy and enhanced robustness compared to DPS across all considered applications, and allows for a reduction in the required number of denoising steps compared to DPS. Our results for the linear conditioning case compare well with those sampled with the preconditioned Crank–Nicolson (pCN) MCMC algorithm (Cotter et al., 2013). The CDPS can easily be adapted to different DM formulations and data error levels. We further present the algorithm for DDPM and DDIM, with example applications, in the Appendix.
@@ 2.1 Diffusion modeling
Diffusion models enable deep generative modeling by transforming samples from a tractable distribution (often Gaussian) into samples from a formally unknown target pdf $p(\textbf{x}_{0})$ on $\mathbb{R}^{d}$ from which samples are available (the TI). The diffusion process typically involves corrupting these samples, $\textbf{x}_{0}\sim p(\textbf{x}_{0})$, through the progressive addition of Gaussian noise at different time-steps $t$, thereby defining the forward process $p(\textbf{x}_{t}|\textbf{x}_{t-1})$. After a sufficient time $T$, such corrupted data distribution $p(\textbf{x}_{T})$ can be considered Gaussian (Ho et al., 2020; Sohl-Dickstein et al., 2015). The backward process, $q(\textbf{x}_{0}|\textbf{x}_{T:1})$, represents the sampling or generative process, which is analytically intractable for an arbitrary $p(\textbf{x}_{0})$. Numerical methods can be used instead, by training a deep neural network $S_{\theta}(-)$ (Ho et al., 2020; Sohl-Dickstein et al., 2015).
@@
where $\nabla_{\textbf{x}_{t}}\log p_{t}(\textbf{x}_{t})$ is the score function of the noisy data distribution $\textbf{x}_{t}$. The generation of samples using Eq. 2 is possible using Langevin dynamics, implying that a distribution of samples $p(\textbf{x}_{0}|\textbf{x}_{T})$ can be obtained given a single $\textbf{x}_{T}$. As an alternative, deterministic sampling can be performed using the ordinary differential equation (ODE) (Song et al., 2021b)
@@
where $\sigma_{t}$ represents the time-dependent noise magnitude.Using Tweedie’s formula (Robbins, 1956; Efron, 2011) the score function is
@@
$\textbf{x}_{t-1}=\sqrt{\alpha_{t-1}}\left(\frac{\textbf{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\ \epsilon_{\theta}^{(t)}}{\sqrt{\bar{\alpha}_{t}}}\right)+\sqrt{1-\bar{\alpha}_{t-1}-\tilde{\sigma}_{t}^{2}}\epsilon_{\theta}^{(t)}+\tilde{\sigma}_{t}\epsilon_{t},$ (8)
where $\beta_{t}$ defines the noise schedule of the diffusion process, $\alpha_{t}=1-\beta_{t}$, $\hat{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$ and $\tilde{\sigma}_{t}^{2}$ is the conditional variance sequence (injected noise during sampling), and $\epsilon_{t}\sim\mathcal{N}(0,I)$. DDIM and DDPM are discretizations of score-based diffusion in continuous-time domain, with $\epsilon^{(t)}=\nabla_{\textbf{x}_{t}}\log p_{t}(\textbf{x}_{t})/\sqrt{1-\bar{\alpha}_{t}}$ (Song et al., 2021b; Dhariwal and Nichol, 2021). In DDPM $\tilde{\sigma}_{t}=\sqrt{\frac{1-\bar{\alpha_{t}}}{1-\alpha_{t}}}\sqrt{1-\frac{\alpha_{t}}{\alpha_{t-1}}}$ and the sampling process is associated to a VP SDE (Ho et al., 2020; Song et al., 2021b). In DDIM, $\tilde{\sigma}_{t}$ can be arbitrarily changed to 00 so that Eq. 8 becomes a VE ODE for deterministic sampling (Song et al., 2021a; Karras et al., 2022).

### v2

@@ Abstract
Diffusion models offer stable training and state-of-the-art performance for deep generative modeling tasks. Here, we consider their use in the context of multivariate subsurface modeling and probabilistic inversion. We first demonstrate that diffusion models enhance multivariate modeling capabilities compared to variational autoencoders and generative adversarial networks. In diffusion modeling, the generative process involves a comparatively large number of generative steps with update rules that can be modified to account for conditioning data. We propose different corrections to the popular Diffusion Posterior Sampling approach by Chung et al. (2023). In particular, we introduce a likelihood approximation accounting for the noise-contamination that is inherent in diffusion modeling. We assess performance in a multivariate geological scenario involving facies and correlated acoustic impedances. Conditional modeling is demonstrated using both local hard data (well logs) and nonlinear geophysics (fullstack seismic data). Our tests show significantly improved statistical robustness, enhanced sampling of the posterior probability density function and reduced computational costs, compared to the original approach. The method can be used with both hard and indirect conditioning data, individually or simultaneously. As the inversion is included within the diffusion process, it is faster than other methods requiring an outer-loop around the generative model, such as Markov chain Monte Carlo.
@@ 1 Introduction
Predicting the spatial distribution of subsurface properties is a key challenge in the Earth and environmental sciences. Probabilistic inverse modeling (Grana et al., 2021; Rubin and Hubbard, 2005) aims to approximate the posterior probability density function (pdf) by integrating prior pdfs with observed data in the form of a likelihood function (Tarantola, 2005). Conditioning data may consist of local measurements (e.g., well logs) or indirect geophysical data describing averaged subsurface properties through linear or nonlinear physical forward operators. Prior pdfs expressed as explicit statistical models (e.g., two-point geostatistical models) are computationally convenient, yet are often gross simplifications of geological heterogeneity, leading to biased predictions and incomplete uncertainty quantification (e.g., Linde et al., 2017; Gómez-Hernández and Wen, 1998; Neuweiler et al., 2011; Journel and Deutsch, 1993; Zhou et al., 2018). Alternatively, prior geological knowledge can be represented using 2D or 3D gridded data, defined as training images (TI) that implicitly describe lower- and higher-order spatial relationships. Multiple-Point Statistics (MPS) simulation methods (Guardiano and Srivastava, 1993; Strebelle, 2002; Mariethoz et al., 2010a; Mariethoz and Caers, 2014) rely on TIs to draw geologically-realistic model realizations and have found wide applicability in both hydro-geology (e.g., Høyer et al., 2017; Le Coz et al., 2017; Straubhaar et al., 2020) and reservoir characterization (e.g., Melnikova et al., 2015; Caers and Zhang, 2004, & references therein). These realizations can easily be made to honor local observations (Straubhaar et al., 2016; Hansen et al., 2018; Straubhaar and Renard, 2021), but struggle with high computational costs and poor posterior exploration when used in the context of nonlinear inversion (e.g., Zahner et al., 2016; Levy et al., 2024; Laloy et al., 2016; Hansen et al., 2012; Mariethoz et al., 2010b).
@@
The substantial progress of deep generative models over the past decade (Goodfellow et al., 2016) have created a strong interest in leveraging such techniques in the geosciences. Initial work considered convolution-based deep neural networks, such as variational autoencoders (VAE; Kingma and Welling, 2013) and generative adversarial networks (GAN; Goodfellow et al., 2014). These methods can encode complex geological patterns of a TI within a low-dimensional manifold, defining prior pdfs (Laloy et al., 2017, 2018). Inversion can then be performed within the latent space of the pre-trained networks (Laloy et al., 2017, 2018; Oliviero-Durmus et al., 2025) using sampling methods, such as Markov chain Monte Carlo (MCMC; Laloy et al., 2017, 2018; Levy et al., 2021; Mosser et al., 2020; Liu et al., 2022) and Sequential Monte Carlo (SMC; Doucet et al., 2001; Amaya et al., 2021), ensemble-based approaches (Canchumuni et al., 2019; Liu and Grana, 2020; Mo et al., 2020), deterministic optimization (Laloy et al., 2019; Dupont et al., 2018), or variational inference (Chan and Elsheikh, 2019; Levy et al., 2023; Miele et al., 2024). Contrarily to MPS algorithms, conditioning to local data remains cumbersome. One can train the generative models as inference methods on specific conditional data, but this necessitates retraining for each application task (e.g. Mosser et al., 2018; Zhang et al., 2021; Laloy et al., 2020; Feng et al., 2024; Miele and Azevedo, 2024). Moreover, GANs are prone to unstable training and mode collapse (e.g., Lucic et al., 2018), and the non-linearity of the generative process often adversely affects the inversion results (Lopez-Alvis et al., 2021; Laloy et al., 2019). Formulations based on VAEs are more stable, but offer rather lossy representations (e.g., Levy et al., 2023; Miele et al., 2024).
Diffusion models (DM; Ho et al., 2020; Sohl-Dickstein et al., 2015) represent the current state-of-the-art in deep generative modeling, outperforming GANs in terms of accuracy without being affected by mode collapse or unstable training (e.g., Dhariwal and Nichol, 2021). These methods are based on reversing a data corruption process transforming samples of a target distribution (the TI) into those of a known distribution (typically Gaussian). The generative process involves iteratively denoising samples originating from the latter, with the denoising trajectory being learned by a deep neural network (e.g., Song et al., 2021a; Ho et al., 2020; Song et al., 2021b; Karras et al., 2022).
In geological subsurface modeling, diffusion models (DMs) have been explored in different contexts. Mosser (2023) showed that autoregressive DMs (Hoogeboom et al., 2022) share structural similarities with sequential geostatistical modeling and demonstrated their potential for simulating multi-scale 2D synthetic channel patterns. Xu et al. (2024) found high modeling accuracy of DMs in analogous geological scenarios and proposed a conditional generative approach, where sparse point observations were incorporated as input channels during training. Aouf et al. (2025) demonstrated that DMs can be used for 2D and 3D modeling of complex porous geological media, further performing conditional modeling on total porosity values, by training the DM to embed such conditioning parameter and associate it to the desired output. Conditioning on embedded well-log data was further investigated by Lee et al. (2025) for latent diffusion models (LDM; Rombach et al., 2022), which perform diffusion within the latent space of a pre-trained encoder-decoder network. Ovanger et al. (2025) extended this work by providing a thorough geostatistical assessment of LDMs against truncated Gaussian random field modeling (Matheron et al., 1987; Mannseth, 2014) for multi-facies scenarios. The authors present overall high modeling performances, but note biases due to the LDM data compression. Diffusion models were also recently introduced to solve inverse modeling problems in the Earth sciences. Wang et al. (2023) used pre-trained DMs as regularizers within a deterministic full waveform inversion (FWI) of seismic data, integrating the FWI’s iterative updates with DM sampling. This framework, designed for subsurface wavespeed predictions, was extended to multiple elastic properties by Taufik et al. (2024). Both studies demonstrate improved accuracy and more realistic results compared to conventional FWI for synthetic geological scenarios. Wang et al. (2024) further proposed a controllable DM-based FWI, where the model learns the relationship between embedded conditioning data and wavespeed distributions. Finally, Di Federico and Durlofsky (2025) show that stochastic inverse modeling can be carried out by combining a pre-trained unconditioned LDM with an ensemble-based inversion method; the authors demonstrated their approach on facies prediction tasks in the context of history matching.
@@
Beyond Earth sciences, the field of inverse modeling with DMs is of broad interest for computer vision and biomedical imaging, and several methodologies were recently proposed; we refer to Daras et al. (2024) and Zhao et al. (2025) for comprehensive overviews. Analogously to methods based on GANs and VAEs, direct training on conditioning data is a popular approach (e.g., Ho and Salimans, 2021; Dhariwal and Nichol, 2021; Whang et al., 2022; Liu et al., 2023; Xu et al., 2024; Lee et al., 2025; Aouf et al., 2025; Wang et al., 2024), but implies computationally demanding re-training when considering new data types or designs. Alternatively, one can rely on inference frameworks that iterate outside the generative process of pre-trained unconditional DMs. Conventional sampling algorithms (e.g., MCMC) are often inefficient in this setting, due to the DM’s large dimensionality and comparatively long generation times (Zhao et al., 2025). More efficient variational inference algorithms have been proposed for posterior approximation with surrogate modeling, but they often struggle to capture complex or multimodal posterior distributions (e.g., Mardani et al., 2023; Alkan et al., 2023; Feng and Bouman, 2024). Ensemble-based approaches (e.g., Di Federico and Durlofsky, 2025) can be limited by comparatively high computational costs; LDMs can be used to reduce dimensionality, but the associated encoder-decoder networks (e.g., VAE or GAN) introduce additional complications such as reduced modeling ability (e.g., leading to bias or smoothing; Ovanger et al., 2025) and non-linearities (Daras et al., 2024).
@@
Another family of conditioning approaches leverage the sequential nature of the denoising process. For example, SMC can be used to progressively sample, at different denoising steps, a posterior distribution that evolves towards the posterior pdf (e.g., Trippe et al., 2022; Cardoso et al., 2023; Wu et al., 2024; Dou and Song, 2024). These asymptotically exact methods work under the assumption of infinite sampling and linear physics. As an alternative, the denoising process itself can be adapted to achieve approximate Bayesian posterior sampling at low computational costs (Song et al., 2021b), provided that the intractable likelihood is approximated at each denoising step. The approximated likelihood score is then combined with the pre-trained denoising score to steer the denoising generative process towards posterior samples. Among numerous such approaches (see e.g., Daras et al., 2024), Diffusion Posterior Sampling (DPS) (Chung et al., 2023) has gained wide popularity as it allows inference for both linear and nonlinear inverse problems. Nonetheless, this class of conditional samplers are affected by the likelihood approximation (Daras et al., 2024). For instance, the accuracy of DPS decreases with the conditioning data noise magnitude and increasing nonlinearity of the problem at hand (Chung et al., 2023). Various adjustments have been proposed in the implementation of DPS, including ad-hoc weighting of the likelihood score. Several studies have highlighted these limitations and proposed partial improvements for linear inverse problems (e.g., Hamidi and Yang, 2025; Boys et al., 2024; Yismaw et al., 2025; Song et al., 2023).
@@
In this work, we investigate DMs for multivariate subsurface geological modeling and propose improvements to DPS for inversion tasks. Our corrected DPS (CDPS) relies on a likelihood approximation that considers not only the observational error of the conditioning data, but also the impact of the noise contamination at the different stages of the DM. The modified likelihood approximation removes the need to arbitrarily weight the likelihood score and is expected to improve robustness and uncertainty quantification. We further highlight inconsistencies in the original implementation of DPS by Chung et al. (2023) implemented with Denoising Diffusion Probabilistic and Implicit Modeling (DDPM and DDIM; Ho et al., 2020; Song et al., 2021a) and suggest corrections. In our study, we rely on a state-of-the-art DM implementation (EDM; Karras et al., 2022) to model a bivariate prior pdf describing sedimentary channel sequences and correlated acoustic impedances. We first demonstrate substantial increases in prior modeling performance against a GAN and a VAE previously introduced for analogous tasks (Miele et al., 2024). Then, we use CDPS to condition realizations on both local observations (well logs) and nonlinear geophysical data (fullstack seismic data), considering Gaussian data noise of different magnitudes. We demonstrate that CDPS yields significantly improved accuracy and enhanced robustness compared to DPS across all considered applications, while allowing for a reduction in the required number of denoising steps. Our results for the linear conditioning case compare well with those sampled with the preconditioned Crank–Nicolson (pCN) MCMC algorithm (Cotter et al., 2013). The CDPS can easily be adapted to different DM formulations and data error levels. We further present the algorithm for DDPM and DDIM, with example applications, in the Appendix.
@@ 2.1 Diffusion modeling
Diffusion models enable deep generative modeling by transforming samples from a tractable distribution (often Gaussian) into samples from an unknown target pdf $p(\textbf{x}_{0})$ on $\mathbb{R}^{d}$, from which samples are available (the TI). The diffusion process typically involves corrupting these samples, $\textbf{x}_{0}\sim p(\textbf{x}_{0})$, through the progressive addition of Gaussian noise at different time-steps $t$, thereby defining the forward process $p(\textbf{x}_{t}|\textbf{x}_{t-1})$. After a sufficient time $T$, the corrupted data distribution $p(\textbf{x}_{T})$ can be considered Gaussian (Ho et al., 2020; Sohl-Dickstein et al., 2015). The backward process, $q(\textbf{x}_{0}|\textbf{x}_{T:1})$, represents the sampling or generative process, which is analytically intractable for an arbitrary $p(\textbf{x}_{0})$. Numerical methods can be used instead by relying on a trained deep neural network $S_{\theta}(-)$ (Ho et al., 2020; Sohl-Dickstein et al., 2015).
@@
where $\nabla_{\textbf{x}_{t}}\log p_{t}(\textbf{x}_{t})$ is the score function of the noisy data distribution $\textbf{x}_{t}$. The generation of samples using Eq. 2 is possible using Langevin dynamics, implying that a distribution of samples $p(\textbf{x}_{0}|\textbf{x}_{T})$ can be obtained from a single $\textbf{x}_{T}$. As an alternative, deterministic sampling can be performed using the ordinary differential equation (ODE) (Song et al., 2021b)
@@
where $\sigma_{t}$ represents the time-dependent noise magnitude. Using Tweedie’s formula (Robbins, 1956; Efron, 2011), the score function is
@@
$\textbf{x}_{t-1}=\sqrt{\bar{\alpha}_{t-1}}\left(\frac{\textbf{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\ \epsilon_{\theta}^{(t)}}{\sqrt{\bar{\alpha}_{t}}}\right)+\sqrt{1-\bar{\alpha}_{t-1}-\tilde{\sigma}_{t}^{2}}\epsilon_{\theta}^{(t)}+\tilde{\sigma}_{t}\epsilon_{t},$ (8)
where $\beta_{t}$ defines the noise schedule of the diffusion process, $\alpha_{t}=1-\beta_{t}$, $\hat{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$ and $\tilde{\sigma}_{t}^{2}$ is the conditional variance sequence (injected noise during sampling), and $\epsilon_{t}\sim\mathcal{N}(0,I)$. DDIM and DDPM are discretizations of score-based diffusion, with $\epsilon^{(t)}=-\nabla_{\textbf{x}_{t}}\log p_{t}(\textbf{x}_{t})\sqrt{1-\bar{\alpha}_{t}}$ (Song et al., 2021b; Dhariwal and Nichol, 2021). In DDPM, $\tilde{\sigma}_{t}=\sqrt{\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}}\sqrt{\frac{1-\bar{\alpha}_{t}}{\bar{\alpha}_{t-1}}}$, while in the DDIM framework $\tilde{\sigma}_{t}$ is set to 0 for deterministic sampling (Ho et al., 2020; Song et al., 2021b), and can hence be associated to a VP SDE and VP ODE, respectively.